Assume that a car decelerates at 2.0m/s2 and comes to a stop after traveling 25m.


a) Find the speed of the car at the start of the deceleration and


b) Find the time required to come to a stop.


1
Expert's answer
2022-03-06T18:05:35-0500

"v=\sqrt{2as}=10~\frac ms,"

"t=\frac va=5~s."